A radiator contains 15 quarts of fluid, 10% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 30% antifreeze? (Round your answer to two decimal places.)
Answer by richwmiller(17219)

0.1*15-0.1x+1x=0.3(15)
1.5+0.9x=4.5
0.9x=4.5-1.5
0.9x=3
x=3.33333333
check
0.1*11.6666667+1*3.33333333=0.3(15)
1.16666667+3.33333333=4.5
4.5=4.5
ok
codemixrep